OpenStudy (anonymous):

If the sum of two numbers x and y is 12, what is the maximum product of (x^3)(y)?

OpenStudy (inkyvoyd):

x+y=12 seek to optimize (x^3)(y) remember that =12-x

OpenStudy (inkyvoyd):

I mean, y=12-x, sorry.

OpenStudy (anonymous):

x^3(12-x), what's the next step? How do I find x and y?

OpenStudy (inkyvoyd):

Remember: you seek to maximize p(x)x^3(12-x) thus you search for where p'(x)=0

OpenStudy (inkyvoyd):

Well, not necessarily always p'(x)=0, but in thsi case you seek it.

OpenStudy (inkyvoyd):

correction: p(x)=x^3(12-x) my apologies openstudy is lagging today and not responding well to my keyboard.

OpenStudy (anonymous):

y = 12-x. f(x) = (x^3)y = (x^3)(12-x). Find the first derivative of f(x). try to find out where x has the absolute maximum. then once u find that x, plug it into y=-x+12 to find y

OpenStudy (raden):

looks the max value will be 3^7

OpenStudy (inkyvoyd):

@RadEn , both of us were trying to not blurt out the answer :/